Test the conditional convergence of the series


1-1/root2+1/root3-1/root4+....
1
Expert's answer
2020-03-10T10:38:41-0400

"1-\frac{1}{\sqrt2}+\frac{1}{\sqrt3}-\frac{1}{\sqrt4}+...+\frac{(-1)^{n-1}}{\sqrt n}+...(*)\\
a_n=\frac{(-1)^{n-1}}{\sqrt n}=\frac{(-1)^{n-1}}{ n^\frac{1}{2}}."

Consider a series of modules

"1+\frac{1}{\sqrt2}+\frac{1}{\sqrt3}+\frac{1}{\sqrt4}+...+\frac{1}{\sqrt n}+...(**)\\"

Series

"\sum\limits_{n=1}^{\infty} \frac{1}{n^\alpha}"

convergent if "\alpha>1" , divergent if "\alpha \leq1" .

In our case "\alpha=\frac{1}{2}<1" then series (**) is divergent.

We examine the series (*) on the basis of Leibniz

"1) |a_1|>|a_2|>...>|a_n|>..."

in our case

"1>\frac{1}{\sqrt 2}>\frac{1}{\sqrt 3}>...>\frac{1}{\sqrt n}>...\\
2) \lim\limits_{n\mapsto \infty}|a_n|=\lim\limits_{n\mapsto \infty} \frac{1}{\sqrt n}=0."

According to Leibniz, the series (*) is convergent.

 Since the series (**) is divergent and the series (*) is convergent, the series (*) is conventionally convergent.